# Salter’s question on the image of the Burau representation of $B_{4}$

### v1

@@ 1 Introduction
In this paper, assuming that $\beta_{4,\,u}$ (resp. $\beta_{4,\,r}$) is faithful, we prove that $\overline{\beta_{4,\,u}}$ (resp. $\overline{\beta_{4,\,r}}$) is not surjective. This is achieved by constructing two types of counterexamples: the first (Theorem 1.1) is hereditary from those found in [20], while the second (Theorem 1.2) is newly discovered.
@@
The representation $\beta_{4,\,r}|_{C_{B_{4}}\left(\sigma_{3}\right)}$ is faithful. Moreover, this faithfulness is preserved modulo $p$ for every prime $p$. In particular, the matrices $S$ and $T^{-1}ST$ generate a free group of rank 2 modulo $p$ for every $p$.
@@
Several authors [10, 11, 31] have studied the two-dimensional Bruhat-Tits building on which the Burau image acts, relating the faithfulness problem to the geometry of buildings. This connection will be explored in the final part of the paper. We examine a large Bruhat-Tits building on which the unitary group, defined with $\mathbb{C}\left[t,\,t^{-1}\right]$ entries, acts. We show that the unipotent kernel $\mathcal{U}_{\mathbb{C}}$ (see Definition 4.4) acts simply transitively on the vertices of this building. Under these conditions, the group $\mathcal{U}_{\mathbb{C}}$ admits a triangle presentation as a locally infinite $\widetilde{A_{2}}$ group (see [7, 8]). More generally, we establish the following result.
@@
Finally, we examine a locally finite connected simplicial subcomplex $\Delta_{\mathbb{C}}\left(\mathcal{H}\right)$ of the building, which is spanned by a finitely generated subgroup $\mathcal{H}$ of $\mathcal{U}_{\mathbb{C}}$ which contains a finite index subgroup of $\left\langle S,\,T\right\rangle$ up to some conjugation (see Section 4 for definitions). An analysis of this subcomplex leads to the following criterion for faithfulness.
@@ 3 Counterexamples in the Stabilizer
Gorin–Lin [15] pointed out that these two generate a free normal subgroup of $B_{4}$, and the quotient map $\rho:B_{4}\to\left\langle\sigma_{2},\,\sigma_{3}\right\rangle$ includes a split short exact sequence, so that $B_{4}$ is a semidirect product $\left\langle b_{1},\,b_{2}\right\rangle\rtimes\left\langle\sigma_{2},\,\sigma_{3}\right\rangle$. Indeed, the following computation is direct from the braid relations of $B_{4}$.
@@
From the equation (3) and (4), we have the following first two elements other than $b_{1}$.
@@
Combining (5) and (6), we obtain the following relation:
$\left(\sigma_{3}\sigma_{2}\sigma_{3}\right)^{2}b_{1}\left(\sigma_{3}\sigma_{2}\sigma_{3}\right)^{-2}$ $=b_{1}^{-1}\left(\left(\sigma_{3}\sigma_{2}\sigma_{3}\right)^{-2}b_{1}\left(\sigma_{3}\sigma_{2}\sigma_{3}\right)^{2}\right)b_{1},$ (7)
which allows us to inductively express all other generators as an appropriate product of $b_{1}$ and $b_{2}b_{1}b_{2}^{-1}$.
@@
Let us momentarily set aside the counterexamples and focus on the faithfulness problem in the centralizer. For the usual Burau representation $\beta_{4,\,r}$, since $\beta_{4,\,r}$ commutes with the Gorin–Lin quotient $\rho$ ([5]), it suffices to show that $\left\langle S^{\prime},\,T^{\prime}S^{\prime}T^{\prime-1}\right\rangle$ is free of rank 2. This result was established by Moran in [26], but he used the order on the real field $\mathbb{R}$, which creates challenges when extending the result to the mod $p$ Burau, where the coefficient field has positive characteristic. Moreover, in general, the Gorin–Lin quotient $\rho$ does not necessarily commute with $\beta_{4,\,r}$ modulo $p$ (for instance, see the counterexamples in [10]).
@@
We may also consider the mod $p$ (or even mod $n$) tame subgroup $\mathcal{T}_{p}$ and the mod $p$ stable subgroup $\mathcal{S}_{p}$ as subgroups of $B_{4}$. By definition, the group $\mathcal{T}_{p}$ (resp. $\mathcal{S}_{p}$) contains $\mathcal{T}$ (resp. $\mathcal{S}_{p}$). In the light of Lemma 3.9 (1), the quotient $\mathcal{T}_{p}/\left\langle b_{1},\,b_{2}\right\rangle$ is isomorphic to the (upper) Hecke congruence subgroup modulo the center, as expected by [21].
@@ 4 The Entire Unitary Group
A natural approach would be to act this unitary group on the associated Bruhat-Tits building. For general definitions of buildings, see [1, 31, 10]. In brief, for a field $\mathbb{F}$, we consider the two-dimensional Bruhat-Tits building $X_{\mathbb{F}\left(t\right)}$ with the valuation at infinity $v_{\infty}$, which is a simplicial complex whose vertices are $3\times 3$ lattices identified along with a uniformizing element $\pi=t^{-1}$. We consider the action by the left multiplication $\mathrm{PGL}\left(3,\,\mathbb{F}\left[t,\,t^{-1}\right]\right)$ on $X_{\mathbb{F}\left(t\right)}$.
@@
Finally, for each orthogonal matrix $A\in\mathrm{O}\left(2,\,\mathbb{E}\right)$, the orthogonal generator $o_{\mathbb{E}}\left[A\right]$ are defined to be
@@
where for the last equality, we used the fact $1+r^{4}\neq 0$.
@@
where $r_{1},\,r_{2}>0$ and $r_{1}r_{2}>1$; and $\left[k_{\mathbb{E}}\left[r\right],\,o_{\mathbb{E}}\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right]\right]=1$ for $r\geq 0$ or $r=\infty$.
To proceed further, we ask whether there is a subgroup of $\mathrm{PU}\left(\mathbb{E}\right)$ that acts transitively on the building with small stabilizers. At this stage, it is challenging to find a suitable answer for the fields $\mathbb{R}$ or $\mathbb{E}$. Let us extend the field once more. For a Euclidean field $\mathbb{E}$, we consider the field $\mathbb{E}\left(i\right)$, where $i^{2}=-1$. An important distinction arises here: the bar involution $\overline{\left(\right)}$ we defined in Section 1 for the Laurent polynomial ring $\mathbb{E}\left[t,\,t^{-1}\right]$ must now be extended compatibly with $i$. We extend the involution in $\mathbb{E}\left(i\right)\left[t,\,t^{-1}\right]$ such that $\overline{i}=-i$. Then, we construct the building $X_{\mathbb{E}\left(i\right)\left(t\right)}$ in the same manner and define the projective unitary group $\mathrm{PU}\left(\mathbb{E}\left(i\right)\right)$.
@@
$B|_{t=i}$ $=\left[\begin{array}[]{ccc}1&*&*\\ 0&1&*\\ 0&0&1\end{array}\right].$
@@
For a Euclidean field $\mathbb{E}$, the unipotent kernel $\mathcal{U}_{\mathbb{E}\left(i\right)}$ is generated by the unipotent generators and their conjugates by unitary generators $u_{\mathbb{E}}\left[A\right]$ such that $A\in\mathrm{SU}\left(2,\,\mathbb{E}\left(i\right)\right)$. The group $\mathcal{U}_{\mathbb{E}\left(i\right)}$ acts on the building $X_{\mathbb{E}\left(i\right)\left(t\right)}$ transitively with the trivial stabilizers, and is a normal subgroup of $\mathrm{PU}\left(\mathbb{E}\left(i\right)\right)$.
@@
Thus, the unipotent kernel $\mathcal{U}_{\mathbb{E}\left(i\right)}$ includes $\mathbf{k}_{\mathbb{E}\left(i\right)}\left[r\right]$ and its every conjugate by unitary generators $u_{\mathbb{E}}\left[A\right]$ such that $A\in\mathrm{SU}\left(2,\,\mathbb{E}\left(i\right)\right)$. As in the proof of Lemma 4.3, these elements represent the type 1 lattices in $\mathrm{Lk}_{\left[I\right]}$, which implies that $\mathcal{U}_{\mathbb{E}\left(i\right)}$ acts transitively on $X_{\mathbb{E}\left(i\right)\left(t\right)}$. Moreover, the stabilizer of the identity lattice $\left[I\right]$ is trivial, since the only unipotent matrix in $\mathrm{U}\left(2,\,\mathbb{E}\left(i\right)\right)$ is the identity matrix.
@@
$u_{\mathbb{E}\left(i\right)}\left[\begin{array}[]{cc}a+ib&c+id\\ -c+id&a-ib\end{array}\right]B|_{t=-i}u_{\mathbb{E}\left(i\right)}\left[\begin{array}[]{cc}a+ib&c+id\\ -c+id&a-ib\end{array}\right]^{-1}$
$=$ $\left[\begin{array}[]{ccc}1&0&0\\ 0&a+ib&c+id\\ 0&-c+id&a-ib\end{array}\right]\left[\begin{array}[]{ccc}1&b_{12}&b_{13}\\ 0&1&0\\ 0&0&1\end{array}\right]\left[\begin{array}[]{ccc}1&0&0\\ 0&a-ib&-c-id\\ 0&c-id&a+ib\end{array}\right]$
$=$ $\left[\begin{array}[]{ccc}1&\left(a-ib\right)b_{12}+\left(c-id\right)b_{13}&-\left(c+id\right)b_{12}+\left(a+ib\right)b_{13}\\ 0&1&0\\ 0&0&1\end{array}\right].$
Since the whole group $\mathrm{PU}\left(\mathbb{E}\left(i\right)\right)$ is generated by the unipotent kernel and the unitary generators, it concludes the proof.
@@
All the arguments presented so far apply to general Euclidean fields. Therefore, it suffices to show that for any given infinite cardinality $\kappa$, there exists a Euclidean field $\mathbb{E}$ of cardinality $\kappa$. Since the theory of ordered fields is (countably) first-order, the theory of Euclidean fields is also first-order. There are at least two models of Euclidean fields, the field of real constructible numbers, and the field of complex numbers $\mathbb{R}$. By applying the upward Loewenheim–Skolem theorem [23], we conclude the proof.
@@
$g_{3}$ $:=u_{\mathbb{C}}\left[\begin{array}[]{cc}i&0\\ 0&1\end{array}\right]\mathbf{k}_{\mathbb{C}}\left[1\right]u_{\mathbb{C}}\left[\begin{array}[]{cc}-i&0\\ 0&1\end{array}\right],\,g_{4}:=u_{\mathbb{C}}\left[\begin{array}[]{cc}0&i\\ i&0\end{array}\right]\mathbf{k}_{\mathbb{C}}\left[1\right]u_{\mathbb{C}}\left[\begin{array}[]{cc}0&-i\\ -i&0\end{array}\right].$
@@
Next, consider the case $u_{\mathbb{E}\left(i\right)}\left[A\right]\mathbf{k}_{\mathbb{E}\left(i\right)}\left[r\right]u_{\mathbb{E}\left(i\right)}\left[A\right]^{-1}$ for a positive element $r\in\mathbb{E}$ and a matrix $A\in\mathrm{SU}\left(2,\,\mathbb{E}\left(i\right)\right)$. Define the map $\phi:\mathcal{U}_{\mathbb{E}\left(i\right)}\to\mathbb{E}\left(i\right)^{2}$ as the composition of the evaluation at $t=-i$ with a map defined by
@@
we have
$\frac{1}{2}$ $\leq\frac{r^{2}}{r^{4}+1},$
or equivalently $r^{2}=1$. Since we have assumed $r$ is positive, we conclude that $r=1$. Then, the right-hand side of (22) reduces to
$\left(\frac{\overline{a_{1}}\left(1+i\right)}{2},\,\frac{-a_{2}\left(1+i\right)}{2}\right),$
@@
We make a brief remark on the computational aspects. The link of the identity lattice within our subcomplex $\Delta_{\mathbb{C}}\left(\mathcal{H}\right)$ is not a spherical building, which implies that $\Delta_{\mathbb{C}}\left(\mathcal{H}\right)$ is not a subbuilding of $X_{\mathbb{E}\left(i\right)\left(t\right)}$. While one might consider the minimal subbuilding $\Delta^{\prime}$ containing $\Delta_{\mathbb{C}}\left(\mathcal{H}\right)$, computations in $\Delta^{\prime}$ are generally quite challenging. This difficulty appears to stem from the properties of the field of real constructible numbers, or at least from the properties of Euclidean fields. Notably, a recent result [24] has established that the field of real constructible numbers is undecidable.

### v2

@@ 1 Introduction
In this paper, assuming that $\beta_{4,\,u}$ (resp. $\beta_{4,\,r}$) is faithful, we prove that $\overline{\beta_{4,\,u}}$ (resp. $\overline{\beta_{4,\,r}}$) is not surjective. This is achieved by constructing two types of counterexamples: the first (Theorem  ${\mathrm{Theorem\,\,1.1}}$) is hereditary from those found in [20], while the second (Theorem 1.2) is newly discovered.
@@
The representation $\beta_{4,\,r}|_{C_{B_{4}}\left(\sigma_{3}\right)}$ is faithful. Moreover, this faithfulness is preserved modulo $p$ for every prime $p$. In particular, the matrices $S$ and $TST^{-1}$ generate a free group of rank 2 modulo $p$ for every $p$.
@@
Several authors [10, 11, 31] have studied the two-dimensional Bruhat-Tits building on which the Burau image acts, relating the faithfulness problem to the geometry of buildings. This connection will be explored in the final part of the paper. We examine a large Bruhat-Tits building $X_{\mathbb{C}\left(t\right)}$ on which the unitary group, defined with $\mathbb{C}\left[t,\,t^{-1}\right]$ entries, acts. We show that the unipotent kernel $\mathcal{U}_{\mathbb{C}}$ (see Definition 4.4) acts simply transitively on the vertices of this building. Under these conditions, the group $\mathcal{U}_{\mathbb{C}}$ admits a triangle presentation as a locally infinite $\widetilde{A_{2}}$ group (see [7, 8]). More generally, we establish the following result.
@@
Finally, we examine a locally finite connected simplicial subcomplex $\Delta_{\mathbb{C}}\left(\mathcal{H}\right)$ of the building, which is spanned by a finitely generated subgroup $\mathcal{H}$ of $\mathcal{U}_{\mathbb{C}}$ which contains a finite index subgroup of $\left\langle S,\,T\right\rangle$ up to some conjugation (see Section 4 for definitions). An analysis of this subcomplex leads to the following criterion for faithfulness. It is worth mentioning that $\Delta_{\mathbb{C}}\left(\mathcal{H}\right)$ is contained in a much smaller building $X_{\mathbb{Q}\left(i\right)\left(t\right)}$.
@@ 3 Counterexamples in the Stabilizer
Gorin–Lin [15] pointed out that these two generate a free normal subgroup of $B_{4}$, and the quotient map $\rho:B_{4}\to\left\langle\sigma_{2},\,\sigma_{3}\right\rangle$ induces a split short exact sequence, so that $B_{4}$ is a semidirect product $\left\langle b_{1},\,b_{2}\right\rangle\rtimes\left\langle\sigma_{2},\,\sigma_{3}\right\rangle$. Indeed, the following computation is direct from the braid relations of $B_{4}$.
@@
From equations (3) and (4), we derive the first two elements, distinct from $b_{1}$, as follows:
@@
By combining equations (5) and (6), we obtain the following relation:
$\left(\sigma_{3}\sigma_{2}\sigma_{3}\right)^{2}b_{1}\left(\sigma_{3}\sigma_{2}\sigma_{3}\right)^{-2}$ $=b_{1}^{-1}\left(\left(\sigma_{3}\sigma_{2}\sigma_{3}\right)^{-2}b_{1}\left(\sigma_{3}\sigma_{2}\sigma_{3}\right)^{2}\right)b_{1}.$ (7)
This relation (7) enables us to inductively express all other generators as an appropriate product of $b_{1}$ and $b_{2}b_{1}b_{2}^{-1}$.
@@
Let us momentarily set aside the counterexamples and focus on the faithfulness problem in the centralizer. For the usual Burau representation $\beta_{4,\,r}$, since $\beta_{4,\,r}$ commutes with the Gorin–Lin quotient $\rho$ [5], it suffices to show that $\left\langle S^{\prime},\,T^{\prime}S^{\prime}T^{\prime-1}\right\rangle$ is free of rank 2. This result was established by Moran in [26], but he used the order on the real field $\mathbb{R}$, which creates challenges when extending the result to the mod $p$ Burau, where the coefficient field has positive characteristic. Moreover, in general, the Gorin–Lin quotient $\rho$ does not necessarily commute with $\beta_{4,\,r}$ modulo $p$ (for instance, see the counterexamples in [10]).
@@
We may also consider the mod $p$ (or even mod $n$) tame subgroup $\mathcal{T}_{p}$ and the mod $p$ stable subgroup $\mathcal{S}_{p}$ as subgroups of $B_{4}$. By definition, the group $\mathcal{T}_{p}$ (resp. $\mathcal{S}_{p}$) contains $\mathcal{T}$ (resp. $\mathcal{S}$). In the light of Lemma 3.9 (1), the quotient $\mathcal{T}_{p}/\left\langle b_{1},\,b_{2}\right\rangle$ is isomorphic to the (upper) Hecke congruence subgroup of level $p$ modulo the center, as expected by [21].
@@ 4 The Entire Unitary Group
A natural approach would be to act this unitary group on the associated Bruhat-Tits building. For general definitions of buildings, see [1, 31, 10]. In brief, for a field $\mathbb{F}$, we consider the two-dimensional Bruhat-Tits building $X_{\mathbb{F}\left(t\right)}$ with the valuation at infinity $v_{\infty}$, which is a simplicial complex whose vertices are $3\times 3$ equivalence classes of lattices along with a uniformizing element $\pi=t^{-1}$. We consider the action by the left multiplication $\mathrm{PGL}\left(3,\,\mathbb{F}\left[t,\,t^{-1}\right]\right)$ on $X_{\mathbb{F}\left(t\right)}$.
@@
Finally, for each orthogonal matrix $A\in\mathrm{O}\left(2,\,\mathbb{E}\right)$, the orthogonal generator $o_{\mathbb{E}}\left[A\right]$ is defined to be
@@
where for the second and the last equalities, we used the fact $01+r^{4}>$ 0.
@@
where $r_{1},\,r_{2}>0$ and $r_{1}r_{2}>1$;
$\left[k_{\mathbb{E}}\left[0\right],\,o_{\mathbb{E}}\left[A\right]\right]$ $=1,$
for each $A\in\mathrm{O}\left(2,\,\mathbb{E}\right)$;
$\left[k_{\mathbb{E}}\left[\infty\right],\,o_{\mathbb{E}}\left[\begin{array}[]{cc}-1&0\\ 0&-1\end{array}\right]\right]$ $=1;$
and
$\left[k_{\mathbb{E}}\left[r\right],\,o_{\mathbb{E}}\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right]\right]$ $=1,$
for $r\geq 0$ or $r=\infty$.
To proceed further, we ask whether there is a subgroup of $\mathrm{PU}\left(\mathbb{E}\right)$ that acts transitively on the vertices of the building with small stabilizers. At this stage, it is challenging to find a suitable answer for the fields $\mathbb{R}$ or $\mathbb{E}$. Let us extend the field once more. For a Euclidean field $\mathbb{E}$, we consider the field $\mathbb{E}\left(i\right)$, where $i^{2}=-1$. An important distinction arises here: the bar involution $\overline{\left(\right)}$ we defined in Section 1 for the Laurent polynomial ring $\mathbb{E}\left[t,\,t^{-1}\right]$ must now be extended compatibly with $i$. We extend the involution in $\mathbb{E}\left(i\right)\left[t,\,t^{-1}\right]$ such that $\overline{i}=-i$. Then, we construct the building $X_{\mathbb{E}\left(i\right)\left(t\right)}$ in the same manner and define the projective unitary group $\mathrm{PU}\left(\mathbb{E}\left(i\right)\right)$.
@@
$B|_{t=-i}$ $=\left[\begin{array}[]{ccc}1&*&*\\ 0&1&*\\ 0&0&1\end{array}\right].$
@@
For a Euclidean field $\mathbb{E}$, the unipotent kernel $\mathcal{U}_{\mathbb{E}\left(i\right)}$ is generated by the unipotent generators and their conjugates by unitary generators $u_{\mathbb{E}}\left[A\right]$ such that $A\in\mathrm{SU}\left(2,\,\mathbb{E}\left(i\right)\right)$. The group $\mathcal{U}_{\mathbb{E}\left(i\right)}$ acts on the building $X_{\mathbb{E}\left(i\right)\left(t\right)}$ transitively on the vertices with the trivial stabilizers, and is a normal subgroup of $\mathrm{PU}\left(\mathbb{E}\left(i\right)\right)$. The quotient map is split, and makes $\mathrm{PU}\left(\mathbb{E}\left(i\right)\right)$ a semidirect product:
$\mathrm{PU}\left(\mathbb{E}\left(i\right)\right)$ $=\mathcal{U}_{\mathbb{E}\left(i\right)}\rtimes\mathrm{U}\left(2,\,\mathbb{E}\left(i\right)\right),$
where the unitary group $\mathrm{U}\left(2,\,\mathbb{E}\left(i\right)\right)$ is identified with $u_{\mathbb{E}\left(i\right)}\left[\mathrm{U}\left(2,\,\mathbb{E}\left(i\right)\right)\right]$.
@@
Thus, the unipotent kernel $\mathcal{U}_{\mathbb{E}\left(i\right)}$ includes $\mathbf{k}_{\mathbb{E}\left(i\right)}\left[r\right]$ and its every conjugate by unitary generators $u_{\mathbb{E}}\left[A\right]$, where $A\in\mathrm{SU}\left(2,\,\mathbb{E}\left(i\right)\right)$. As demonstrated in the proof of Lemma 4.3, these elements represent the type 1 lattices in $\mathrm{Lk}_{\left[I\right]}$. Consequently, $\mathcal{U}_{\mathbb{E}\left(i\right)}$ acts transitively on the vertices of $X_{\mathbb{E}\left(i\right)\left(t\right)}$. Furthermore, the stabilizer of the identity lattice $\left[I\right]$ is trivial, as the identity matrix is the only unipotent element in $\mathrm{U}\left(2,\,\mathbb{E}\left(i\right)\right)$.
@@
Moreover, since $\mathcal{U}_{\mathbb{E}\left(i\right)}$ and $u_{\mathbb{E}\left(i\right)}\left[\mathrm{U}\left(2,\,\mathbb{E}\left(i\right)\right)\right]$ have trivial intersection and together generate $\mathrm{PU}\left(\mathbb{E}\left(i\right)\right)$, the entire group $\mathrm{PU}\left(\mathbb{E}\left(i\right)\right)$ is the semidirect product of these two subgroups.
@@
All the arguments presented so far apply to general Euclidean fields. Therefore, it suffices to show that for any given infinite cardinality $\kappa$, there exists a Euclidean field $\mathbb{E}$ of cardinality $\kappa$. Since the theory of ordered fields is (countably) first-order, the theory of Euclidean fields is also first-order. There are at least two models of Euclidean fields, the field of real constructible numbers, and the field of real numbers $\mathbb{R}$. By applying the upward Loewenheim–Skolem theorem [23], we conclude the proof.
@@
$g_{3}$ $:=u_{\mathbb{C}}\left[\begin{array}[]{cc}i&0\\ 0&-i\end{array}\right]\mathbf{k}_{\mathbb{C}}\left[1\right]u_{\mathbb{C}}\left[\begin{array}[]{cc}-i&0\\ 0&i\end{array}\right],\,g_{4}:=u_{\mathbb{C}}\left[\begin{array}[]{cc}0&i\\ i&0\end{array}\right]\mathbf{k}_{\mathbb{C}}\left[1\right]u_{\mathbb{C}}\left[\begin{array}[]{cc}0&-i\\ -i&0\end{array}\right].$
@@
We establish that the spanning subcomplex $\Delta_{\mathbb{E}\left(i\right)}\left(\mathcal{H}\right)$ is locally finite. In other words, the link of a vertex in $\Delta_{\mathbb{E}\left(i\right)}\left(\mathcal{H}\right)$ contains only finitely many vertices. Notably, the proof of this finiteness relies on a complete (infinite) classification of the generators of $\mathcal{U}_{\mathbb{E}\left(i\right)}$.
@@
Next, consider the case $u_{\mathbb{E}\left(i\right)}\left[A\right]\mathbf{k}_{\mathbb{E}\left(i\right)}\left[r\right]u_{\mathbb{E}\left(i\right)}\left[A\right]^{-1}$ for a nonnegative element $r\in\mathbb{E}$ and a matrix $A\in\mathrm{SU}\left(2,\,\mathbb{E}\left(i\right)\right)$. Define the map $\phi:\mathcal{U}_{\mathbb{E}\left(i\right)}\to\mathbb{E}\left(i\right)^{2}$ as the composition of the evaluation at $t=-i$ with a map defined by
@@
we obtain
$\frac{r^{2}}{r^{4}+1}=0\;\;\mathrm{or}\;\;\frac{1}{2}$ $\leq\frac{r^{2}}{r^{4}+1},$
where the first condition implies $r=0$, which corresponds to $\left[d_{1}^{-1}d_{2}^{-1}\right]$, and the second condition implies $r^{2}=1$. Assuming the latter, we conclude $r=1$, as $r$ is assumed to be nonnegative. Substituting $r=1$, the right-hand side of (22) simplifies to
$\left(\frac{\overline{a_{1}}\left(1+i\right)}{2},\,\frac{-a_{2}\left(1+i\right)}{2}\right).$
@@
We make a brief remark on the computational aspects. The link of the identity lattice within our subcomplex $\Delta_{\mathbb{C}}\left(\mathcal{H}\right)$ is not a spherical building, which implies that $\Delta_{\mathbb{C}}\left(\mathcal{H}\right)$ is not a subbuilding of $X_{\mathbb{C}\left(t\right)}$ (or $X_{\mathbb{Q}\left(i\right)\left(t\right)}$). While one might consider the minimal subbuilding $\Delta^{\prime}$ containing $\Delta_{\mathbb{C}}\left(\mathcal{H}\right)$, computations in $\Delta^{\prime}$ are generally quite challenging. This difficulty appears to stem from the properties of the field of real constructible numbers, or at least from the properties of Euclidean fields. Notably, a recent result [24] has established that the field of real constructible numbers is undecidable.